# GroupMamba: Parameter-Efficient and Accurate Group Visual State Space Model

### v1

@@ 4 Experiments
4.1 Image Classification
Settings: The image classification experiments are based on ImageNet-1K [9], which comprising of over $1.28$ million training images and 50K validation images, spanning $1,000$ categories. Following [35], we train our models for using the AdamW [38] optimizer and a cosine decay learning rate scheduler for $300$ epochs, including a $20$ epoch warm-up. The total batch size is set to $1024$, with models trained on 8x A100 GPUs, each with 80GB of CUDA memory. Optimizer betas are set to $(0.9,0.999)$; momentum is set to $0.9$, and an initial learning rate of $1\times 10^{-3}$ is used with a weight decay of $0.05$. Label smoothing of $0.1$ is used alongside the distillation objective (see Section 3.4).
Results: Table 1 presents a comparison of our proposed GroupMamba models (T, S, B) with various state-of-the-art methods. The GroupMamba models exhibit a notable balance of accuracy and computational efficiency. GroupMamba-T achieves a top-1 accuracy of $83.3\%$ with $23$ million parameters and $4.5$ GFLOPs, outperforming ConvNeXt-T [37] and Swin-T [36] by $1.2\%$ and $2.0\%$, respectively, with fewer parameters. Additionally, GroupMamba-T surpasses the recently introduced SSM models, outperforming VMamba-T [34] and LocalVMamba-T [26] by $0.8\%$ and $0.6\%$, respectively, while using $26\%$ fewer parameters than VMamba-T. GroupMamba-S, with $34$ million parameters and $7.0$ GFLOPs, achieves an accuracy of $83.9\%$, surpassing VMamba-S [34], Swin-S [36], and EfficientVMamba-B [46]. The performance is better than LocalVMamba-S [26] by 0.2% with $32\%$ fewer parameters. Furthermore, GroupMamba-B achieves an accuracy of $84.5\%$ with only $57$ million parameters and $14$ GFLOPs, exceeding VMamba-B [34] by $0.6\%$ while using $36\%$ fewer parameters.
Table 1: Performance comparison of GroupMamba models with state-of-the-art convolution-based, attention-based, and SSM-based models on ImageNet-1K [9]. Our models demonstrate superior performance and achieve a better trade-off between accuracy and model parameters.
@@ 4.2 Object Detection and Instance Segmentation
Settings: We evaluate the performance of GroupMamba-T for object detection on the MS-COCO 2017 dataset [32]. Our method is based on the Mask-RCNN [21] detector with the hyperparameters as used for Swin [36]. We use the AdamW [38] optimizer and train Mask-RCNN with GroupMamba-T backbone for $12$ epochs. The backbone is initialized and fine-tuned from the ImageNet-1K [9]. We use an initial learning rate of $1\times 10^{-4}$ and decay by a factor of $10$ at epochs $9$ and $11$.
Results: Table 2 shows the results of GroupMamba-T, comparing it against various state-of-the-art models for object detection and instance segmentation using the Mask R-CNN framework on the MS-COCO dataset. Our model achieves ${}^{\text{b}}$) of 47.6 and ${}^{\text{m}}$) of 42.9. It surpasses ResNet-50 [22], Swin-T [35], ConvNeXt-T [37]. In addition, GroupMamba-T has competitive performance compared to VMamba-T [34] and LocalVMamba-T [26], with less 20% parameters compared to VMamba-T. In Figure 3, we present the qualitative results of GroupMamba-T on samples from the COCO validation set [32], demonstrating its performance in instance segmentation and object detection. Our model accurately localizes objects and correctly segments them across diverse scenes and varying scales.
@@ 4.3 Semantic Segmentation
Settings: We also evaluate the performance of GroupMamba-T for semantic segmentation on the ADE20K [69] dataset. The framework is based on the UperNet [64] architecture, and we follow the same hyperparameters as used for the Swin [36] backbone. More specifically, we use the AdamW [38] optimizer for a total of $160k$ iterations with an initial learning rate of $6\times 10^{-5}$. The default input resolution used in our experiments is $512\times 512$.
@@
Results: The GroupMamba-T model demonstrates favorable performance in semantic segmentation compared to various state-of-the-art methods, as presented in Table 3. GroupMamba-T achieves a mIoU of $48.6$ in single-scale and $49.2$ in multi-scale evaluation, with $49M$ parameters and $955G$ FLOPs. This outperforms ResNet-50 [22], Swin-T [36], and ConvNeXt-T [37]. Additionally, GroupMamba-T exceeds the performance of the recent SSM methods, including ViM-S [70], VMamba-T [34], and LocalVMamba [26] with fewer number of parameters. In Figure 4, we show qualitative results of GroupMamba-T on samples from the ADE20K [69] validation set for semantic segmentation. The first row shows the ground truth masks, while the second row displays the predicted masks. It is notable that our model delineates the masks accurately, highlighting the effectiveness of semantic segmentation. The quantitative and qualitative results of GroupMamba demonstrate the robust generalization capability of our GroupMamba backbones across diverse downstream tasks, including semantic segmentation, object detection, and instance segmentation.
Table 2: Performance comparison for object detection and instance segmentation on MS-COCO [32] using Mask R-CNN [21]: $AP^{b}$ and $AP^{m}$ signify box AP and mask AP, respectively. FLOPs, are computed for an input dimension of $1280\times 800$.
Mask R-CNN 1$\times$ schedule
[…]
ResNet-50 [22] 38.2 58.8 41.4 34.7 55.7 37.2 44M 260G
Swin-T [36] 42.7 65.2 46.8 39.3 62.2 42.2 48M 267G
ConvNeXt-T [37] 44.2 66.6 48.3 40.1 63.3 42.8 48M 262G
PVTv2-B2 [62] 45.3 67.1 49.6 41.2 64.2 44.4 45M 309G
VMamba-T [34] 47.4 69.5 52.0 42.7 66.3 46.0 50M 270G
LocalVMamba-T [26] 46.7 68.7 50.8 42.2 65.7 45.5 45M 291G
GroupMamba-T 47.6 69.8 52.1 42.9 66.5 46.3 40M 279G
@@
Table 3: Performance comparison for semantic segmentation on ADE20K [69] using UperNet [64]. The terms ’SS’ and ’MS’ refer to evaluations conducted at single-scale and multi-scale levels, respectively. FLOPs are computed for an input dimension of $512\times 2048$.
method crop size mIoU (SS) mIoU (MS) #param. FLOPs
ResNet-50 [22] $512^{2}$ 42.1 42.8 67M 953G
DeiT-S + MLN $512^{2}$ 43.8 45.1 58M 1217G
Swin-T [36] $512^{2}$ 44.4 45.8 60M 945G
ConvNeXt-T [37] $512^{2}$ 46.0 46.7 60M 939G
ViM-S [70] $512^{2}$ 44.9 - 46M -
VMamba-T [34] $512^{2}$ 48.3 48.6 62M 948G
EfficientVMamba-B [46] $512^{2}$ 46.5 47.3 65M 930G
LocalVMamba-T [26] $512^{2}$ 47.9 49.1 57M 970G
GroupMamba-T $512^{2}$ 48.6 49.2 49M 955G
4.4 Ablation Study
Table 4 illustrates the effect of each of our contributions, compared to VMamba-T. The proposed GroupMamba-T without the CAM module achieves a top-1 accuracy of 82.2% in the second row. Integrating the proposed CAM module, as shown in the third row, improves performance by 0.3%, with comparable parameters and FLOPs. In the fourth row, we expand the number of channels to match the parameter count of VMamba-T (without using distillation). Our GroupMamba-T* surpasses VMamba-T by 0.6% with an equivalent number of parameters. In the final row, we incorporate the proposed distillation loss without expanding the channels, which boosts performance by 0.8% compared to VMamba-T while using 26% fewer parameters.
Table 4: Ablation study on GroupMamba-T for the proposed CAM module and Distilled Loss.
Method #Param. FLOPs Top-1 acc.
VMamba-T [34] 31M 4.9G 82.5
GroupMamba-T w/o CAM 22M 4.6G 82.2
GroupMamba-T with CAM 23M 4.6G 82.5
GroupMamba-T* with CAM 31M 5.4G 83.1
GroupMamba-T with CAM + Distilled Loss 23M 4.6G 83.3
5 Conclusion
In this paper, we tackle the computational inefficiencies and stability challenges associated with visual SSMs for computer vision tasks by introducing a novel layer called the Modulated Group Mamba. We also propose a multi-directional scanning method that improves parameter efficiency by scanning in four spatial directions and leveraging the Channel Affinity Modulation (CAM) operator to enhance feature aggregation across channels. To stabilize training, especially for larger models, we employ a distillation-based training objective. Our experimental results demonstrate that the proposed GroupMamba models outperform recent SSMs while requiring fewer parameters.

### v2

@@ 4 Experiments
Table 1: Performance comparison of GroupMamba models with state-of-the-art convolution-based, attention-based, and SSM-based models on ImageNet-1K [9]. Our models demonstrate better trade-off between accuracy and parameters.
4.1 Image Classification
Settings: The image classification experiments are based on ImageNet-1K [9], which comprising of over $1.28$ million training images and 50K validation images, spanning $1,000$ categories. Following [38], we train our models for using the AdamW [39] optimizer and a cosine decay learning rate scheduler for $300$ epochs, including a $20$ epoch warm-up. The total batch size is set to $1024$, with models trained on 8x A100 GPUs, each with 80GB of CUDA memory. Optimizer betas are set to $(0.9,0.999)$; momentum is set to $0.9$, and an initial learning rate of $1\times 10^{-3}$ is used with a weight decay of $0.05$. Label smoothing of $0.1$ is used alongside the distillation objective (see Sec. 3.4).
Detection & Instance Segmentation Semantic Segmentation
[…]
ResNet-50 [21] 38.2 58.8 41.4 34.7 55.7 37.2 44M 260G 42.1 42.8
Swin-T [36] 42.7 65.2 46.8 39.3 62.2 42.2 48M 267G 44.4 45.8
ConvNeXt-T [37] 44.2 66.6 48.3 40.1 63.3 42.8 48M 262G 46.0 46.7
VMamba-T [35] 47.4 69.5 52.0 42.7 66.3 46.0 50M 270G 48.3 48.6
LocalVMamba-T [27] 46.7 68.7 50.8 42.2 65.7 45.5 45M 291G 47.9 49.1
GroupMamba-T 47.6 69.8 52.1 42.9 66.5 46.3 40M 279G 48.6 49.2
Table 2: Comparison of model performance on dense prediction tasks: Object detection and instance segmentation results on MS-COCO [33] using Mask R-CNN 1$\times$ schedule [22], and semantic segmentation results on ADE20K [69] using UperNet [64]. ’SS’ and ’MS’ denote single-scale and multi-scale evaluations, respectively. $AP^{b}$ and $AP^{m}$ represent box and mask AP.
Results: Tab. 1 presents a comparison of our proposed GroupMamba models (T, S, B) with various state-of-the-art methods. The GroupMamba models exhibit a notable balance of accuracy and computational efficiency. GroupMamba-T achieves a top-1 accuracy of $83.3\%$ with $23$ million parameters and $4.5$ GFLOPs, outperforming ConvNeXt-T [37] and Swin-T [36] by $1.2\%$ and $2.0\%$, respectively, with fewer parameters. Additionally, GroupMamba-T surpasses the recently introduced SSM models, outperforming VMamba-T [35] and LocalVMamba-T [27] by $0.8\%$ and $0.6\%$, respectively, while using $26\%$ fewer parameters than VMamba-T. GroupMamba-S, with $34$ million parameters and $7.0$ GFLOPs, achieves an accuracy of $83.9\%$, surpassing VMamba-S [35], Swin-S [36], and EfficientVMamba-B [47]. The performance is better than LocalVMamba-S [27] by 0.2% with $32\%$ fewer parameters. Furthermore, GroupMamba-B achieves an accuracy of $84.5\%$ with only $57$ million parameters and $14$ GFLOPs, exceeding VMamba-B [35] by $0.6\%$ while using $36\%$ fewer parameters.
@@ 4.2 Object Detection and Instance Segmentation
Settings: We evaluate the performance of GroupMamba-T for object detection on the MS-COCO 2017 dataset [33]. Our method is based on the Mask R-CNN 1$\times$ schedule [22] detector with the hyperparameters as used for Swin [36]. We use the AdamW [39] optimizer and train Mask-RCNN with GroupMamba-T backbone for $12$ epochs. The backbone is initialized and fine-tuned from the ImageNet-1K [9]. We use an initial learning rate of $1\times 10^{-4}$ and decay by a factor of $10$ at epochs $9$ and $11$. FLOPs are computed for an input dimension of $1280\times 800$.
Results: Tab. 2 shows the results of GroupMamba-T, comparing it against various state-of-the-art models for object detection and instance segmentation using the Mask R-CNN framework on the MS-COCO dataset. Our model achieves ${}^{\text{b}}$) of 47.6 and ${}^{\text{m}}$) of 42.9. It surpasses ResNet-50 [21], Swin-T [38], ConvNeXt-T [37]. In addition, GroupMamba-T has competitive performance compared to VMamba-T [35] and LocalVMamba-T [27], with less 20% parameters compared to VMamba-T. Fig. 3 (first row) displays qualitative examples of object detection and instance segmentation. GroupMamba-T accurately detects and segments the targets in various scenes. More qualitative examples are presented in the supplementary material.
Figure 3: Qualitative results of GroupMamba-T for object detection and instance segmentation (first row) on the MS-COCO val. set and semantic segmentation (second row) on ADE20k val. set.
@@ 4.3 Semantic Segmentation
Settings: We also evaluate the performance of GroupMamba-T for semantic segmentation on the ADE20K [69] dataset. The framework is based on the UperNet [64] architecture, and we follow the same hyperparameters as used for the Swin [36] backbone. More specifically, we use the AdamW [39] optimizer for a total of $160k$ iterations with an initial learning rate of $6\times 10^{-5}$. The default resolution in our experiments is $512\times 512$.
@@
Results: The GroupMamba-T model demonstrates favorable performance in semantic segmentation compared to various state-of-the-art methods, as presented in Tab. 2. GroupMamba-T achieves a mIoU of $48.6$ in single-scale and $49.2$ in multi-scale evaluation. This outperforms ResNet-50 [21], Swin-T [36], and ConvNeXt-T [37]. Additionally, GroupMamba-T exceeds the performance of the recent SSM methods, including ViM-S [70], VMamba-T [35], and LocalVMamba [27]. Fig. 3 (second row) shows qualitative examples of GroupMamba-T. These examples demonstrate our model’s ability to accurately segment various classes for indoor and outdoor scenes. More qualitative examples are presented in the supplementary material.
Figure 4: Comparison of GroupMamba variants and SSM-based methods in top-1 accuracy on ImageNet-1k [9] and computational efficiency in terms of throughput and number of parameters. The throughput (number of predicted samples per second) is measured using a single NVIDIA A100 GPU with a batch size of 128 for all methods.
4.4 Ablation Study
Fig. 4 shows the impact of each proposed contribution in terms of top-1 accuracy, number of parameters, and throughput, compared to other SSM-based methods. GroupMamba-T with 4-D scanning, comprising 22M parameters, achieves a top-1 accuracy of 82.30% and a throughput of 803. By applying a unidirectional 1D scan across $N/4$ channels in four directions—left-to-right, right-to-left, top-to-bottom, and bottom-to-top instead of the full 4-D scanning across all $N$ channels, the throughput significantly increased from 803 to 1125, with only 0.1% drop in accuracy and the same number of parameters.
The integration of the CAM module further elevates the top-1 accuracy from 82.20% to 82.50%, with a minor reduction in throughput (from 1125 to 1069). Finally, incorporating the proposed distillation-based loss pushes the top-1 accuracy to 83.30%, while preserving the throughput at 1069. Compared to Vim-S [70], GroupMamba-T demonstrates a more efficient design, achieving a 2.8% improvement in top-1 accuracy with 1.5$\times$ higher throughput, all while utilizing fewer parameters. Compared to LocalVMamba-T [27], GroupMamba-T has a 0.6% higher accuracy in top-1 accuracy, with 3$\times$ faster and smaller number of parameters. Regarding VMamba-T V1 [35], our model achieves a 1.1% gain in top-1 accuracy with a comparable number of parameters while being faster by 2.5$\times$. Likewise, when compared to VMamba-T V2 [35], GroupMamba-T shows marginally faster throughput, an increase of 0.8% in top-1 accuracy, and a 26% improvement in parameter efficiency.
5 Conclusion and Future Work
In this paper, we tackle the computational inefficiencies and stability challenges associated with visual SSMs for computer vision tasks by introducing a novel layer called Modulated Group Mamba. We also propose a multi-directional scanning method that improves parameter efficiency by scanning in four spatial directions and leveraging Channel Affinity Modulation (CAM) operator to enhance feature aggregation across channels. To stabilize training, especially for larger models, we employ a distillation-based training objective. Our experiments demonstrate that the proposed GroupMamba models outperform recent SSMs while being more efficient in terms of parameters and throughput.
Our research has focused on image classification, object detection, and segmentation. To further validate and extend the generalization ability of our method, we aim to explore additional downstream tasks, such as video recognition and time-series data applications. Evaluating the Modulated Group Mamba layer in these contexts will help to uncover its potential benefits and limitations, providing deeper insights and guiding further improvements.
6 Acknowledgments
The computations were enabled by resources provided by NAISS at Alvis partially funded by Swedish Research Council through grant agreement no. 2022-06725, LUMI hosted by CSC (Finland) and LUMI consortium, and by Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the NSC.
Syed Talal Wasim and Juergen Gall have been supported by the Federal Ministry of Education and Research (BMBF) under grant no. 01IS22094A WEST-AI and the ERC Consolidator Grant FORHUE (101044724).
Supplementary Material
In this section, we further include more results and analysis to complement the main paper. We provide additional details on the following topics:
  • Architectural Details (Sec. 7)
  • Ablations (Sec. 8)
  • Qualitative Results (Sec. 9)
  • Discussion (Sec. 10)
  • Limitations (Sec. 11)
7 Architectural Details
We develop three variants of our GroupMamba backbones, each tailored to different performance and efficiency requirements: GroupMamba-T (Tiny), GroupMamba-S (Small), and GroupMamba-B (Base), with 23M, 34M, and 57M parameters, respectively. These variants differ in their channel dimensions and the number of layers per stage, as detailed in Tab. 6.
8 Ablations
In Tab. 3, we provide additional ablation results regarding the distillation training objective. For the GroupMamba-T and GroupMamba-S variants, the distilled loss improves performance by an absolute gain of 0.8% and 0.9%, respectively. For the largest variant, GroupMamba-B, the distilled loss improves performance by 1.3%. This demonstrates that larger Mamba-based models with MLP tend to saturate and struggle to converge effectively without distillation. Incorporating distillation for the large model boosts its performance from 83.2% to 84.5%.
We also visualize the training loss curves with and without our proposed distilled loss for GroupMamba-S in Fig. 5. The shaded areas indicate the standard deviation of loss across the training epochs. As shown, incorporating the distilled loss (green curve) consistently leads to lower training losses and less loss variability throughout the training process, leading to improved stability.
We compare in Tab. 4 the performance of different scanning directions with respect to the number of groups for GroupMamba-T. In the first row, we use Direction 1. In the second row, we use Direction 1 and Direction 2. In the last row, we use the four scanning scanning directions (As visualized in Fig. 2 (d). Four groups with four directions capture richer spatial cues, which provide comprehensive feature representation and lead to higher top-1 accuracy with comparable throughput.
We also conduct an ablation study to evaluate efficiency with varying numbers of groups. While utilizing two groups reduces parameters by 15% and four groups achieves a reduction of 26%, employing eight groups yields only a marginally greater reduction of 28% due to the nonlinear scaling of MLP parameters. In addition, using eight groups (with eight scanning directions) decreases throughput, negatively impacting model efficiency. Hence, four groups have the optimal trade-off between parameter reduction and high throughput.
Method #Param. FLOPs Top-1 acc.
GroupMamba-T w/o Distilled Loss 23M 4.6G 82.5
GroupMamba-T with Distilled Loss 23M 4.6G 83.3 (+0.8)
GroupMamba-S w/o Distilled Loss 34M 7.0G 83.0
GroupMamba-S with Distilled Loss 34M 7.0G 83.9 (+0.9)
GroupMamba-B w/o Distilled Loss 57M 14G 83.2
GroupMamba-B with Distilled Loss 57M 14G 84.5 (+1.3)
Table 3: Ablation study on GroupMamba variants with and without the Distilled Loss.
In Tab. 5, we present an additional ablation study and a fair comparison between GroupMamba-T and VMamba-T without distillation alongside another variant of GroupMamba-T designed to match the parameter count of VMamba-T for balanced evaluation. Remarkably, GroupMamba-T achieves equivalent performance to VMamba-T with 26% fewer parameters. When parameter counts are matched, the enhanced variant, GroupMamba-T, outperforms VMamba-T, achieving a top-1 accuracy of 83.1% on ImageNet-1K, compared to 82.5% for VMamba-T, without using any distillation.
Scanning Directions Throughput (ms) $\uparrow$ #Param $\downarrow$ Top-1 (%) $\uparrow$
D1 1096 23M 82.9
D1, D2 1087 23M 83.1
D1, D2, D3, D4 1069 23M 83.3
Table 4: Comparison of different scanning directions in terms of throughput, parameters, and top-1 accuracy for GroupMamba-T.
Figure 5: Training loss visualization for GroupMamba-S with and without the proposed distilled loss.
9 Qualitative Results
In Fig. 6, we show additional qualitative results of GroupMamba-T on samples from the ADE20K [69] validation set for semantic segmentation. The first row shows the ground truth masks, while the second row displays the predicted masks. Our model consistently has sharp and accurate delineations, effectively capturing fine details and complex object boundaries, further emphasizing its robustness in semantic segmentation. Similarly, we present in Fig. 7 additional qualitative results of GroupMamba-T on samples from the COCO validation set [33], showcasing its strong performance in both instance segmentation and object detection tasks. The model excels at accurately localizing objects and producing precise segmentations, even in complex scenes with varying scales, multiple instances, and challenging backgrounds. The quantitative and qualitative results of GroupMamba demonstrate the robust generalization capability of our GroupMamba backbones across diverse downstream tasks, including semantic segmentation, object detection, and instance segmentation.
Method #Param. FLOPs Top-1 acc.
VMamba-T 31M 4.9G 82.50%
GroupMamba-T 23M 4.5G 82.50%
GroupMamba-T 31M 5.2G 83.10%
Table 5: Comparison of VMamba-T and GroupMamba-T without distillation. The number of channels is increased in GroupMamba-T to match the same parameters of VMamba-T
Stage Output Resolution Type Config GroupMamba
T S B
stem $\frac{H}{2}\times\frac{W}{2}$ Patch Embed. Patch Size $k=3\times 3,s=2$
Embed. Dim. 32 64 64
$\frac{H}{4}\times\frac{W}{4}$ Patch Embed. Patch Size $k=3\times 3,s=2$
Embed. Dim. 96 96 128
1 $\frac{H}{4}\times\frac{W}{4}$ Modulated Group Mamba Stage [$C_{1}$, $N_{1}$] 96, 2 96, 2 128, 2
2 $\frac{H}{8}\times\frac{W}{8}$ Down-sampling Patch Size $k=3\times 3,s=2$
Embed. Dim. 192 192 256
$\frac{H}{8}\times\frac{W}{8}$ Modulated Group Mamba Stage [$C_{2}$, $N_{2}$] 192, 2 192, 2 256, 2
3 $\frac{H}{16}\times\frac{W}{16}$ Down-sampling Patch Size $k=3\times 3,s=2$
Embed. Dim. 368 384 496
$\frac{H}{16}\times\frac{W}{16}$ Modulated Group Mamba Stage [$C_{3}$, $N_{3}$] 368, 9 384, 20 496, 20
4 $\frac{H}{32}\times\frac{W}{32}$ Down-sampling Patch Size $k=3\times 3,s=2$
Embed. Dim. 760 768 1012
$\frac{H}{32}\times\frac{W}{32}$ Modulated Group Mamba Stage [$C_{4}$, $N_{4}$] 760, 2 768, 2 1012, 2
Parameters 23M 34M 57M
GFLOPs 4.5G 7.0G 14.0G
Table 6: GroupMamba Architectures. Description of the configurations of each model with respect to the output resolution, the output channels $C$, the number of blocks $N$, and the model’s GFLOPs and parameters. Between two consecutive stages, a down-sampling layer is used to increase the features and reduce the resolution by two.
10 Discussion
Our main contributions include introducing the Modulated Group Mamba layer, which enhances computational efficiency and interaction in state-space models through a multi-direction scanning method. We also introduce the Channel Affinity Modulation (CAM) operator to improve feature aggregation across channels, addressing limitations in grouping operations. Additionally, we employ a distillation-based training objective to stabilize the training of models with a large number of parameters. These contributions enable us to achieve competitive performance with recent state-space models in image classification, object detection, instance segmentation, and semantic segmentation with fewer number of parameters.
This can further facilitate the development of vision foundation models based on Mamba that can be scaled to a large number of parameters efficiently and stably. The Modulated Group Mamba layer and CAM operator enhance computational efficiency and feature interaction, allowing models to manage more extensive and complex datasets without excessive resource demands. The distillation-based training objective ensures stability during training, which is crucial for maintaining performance as model sizes increase. Together, these advancements enable the creation of scalable, reliable vision models that can be deployed effectively in various real-world applications.
@@ 10 Discussion
11 Limitations
Despite demonstrating clear improvements in efficiency, stability, and accuracy for image classification tasks and fewer parameters for dense prediction tasks, our proposed Modulated Group Mamba layer shows relatively comparable performance on downstream tasks such as object detection and segmentation to VMamba. This minor improvement can be attributed to the more complex nature and diverse requirements of these dense prediction tasks, where the accuracy relies heavily not only on effective global dependency capture but also on more localized spatial feature aggregation and specialized detection or segmentation heads. The proposed model architecture enhances parameter efficiency and global feature modeling through SSM mechanisms, but addressing the intricacies inherent to localization-sensitive tasks may require additional targeted modules or task-specific optimizations.
Although the incorporation of knowledge distillation has successfully improved training stability and yielded performance gain for large-scale models, investigating more efficient or self-guided stabilization approaches would help enhance the model training practicality without requiring auxiliary external teacher models.